### v1

@@ Abstract
Keywords:
Pre-trained Models Software Dependencies Model Reuse ML pipeline Software Supply Chain Model Interactions Technical Debt Machine Learning Deep Learning
@@ 1 Introduction
Deep neural networks (DNNs) are driving major advances in domains such as image analysis (Namysl and Konya, 2019; Yin et al., 2019), natural language processing (Collobert and Weston, 2008; Nakano et al., 2021; Zhou, 2022), and speech recognition (Gaikwad et al., 2010; Abdel-Hamid et al., 2014; Nassif et al., 2019). In software engineering, DNNs are increasingly applied to tasks such as code generation and automated testing (Le et al., 2020; Ale and Yarram, 2024; Chen et al., 2025). Yet, training these models from scratch is resource-intensive, requiring substantial computational power and engineering effort as architectures grow in size and complexity (Narayanan et al., 2019; Geiping and Goldstein, 2023). Pre-trained models (PTMs) have emerged as a practical solution to this challenge. By encapsulating both the architecture and the parameters learned from data, PTMs enable developers to transfer knowledge across tasks, significantly reducing training costs and development time (Marcelino, 2018; Han et al., 2021). The recent emergence of foundation models (FMs), i.e., large-scale PTMs trained on broad, heterogeneous data and adaptable to a variety of tasks (Bommasani et al., 2021), has further accelerated PTM reuse, with models such as GPT-4 (OpenAI, 2023a) and Stable Diffusion (Rombach et al., 2022a). Reports indicate that the global AI market, encompassing technologies like PTMs, is projected to grow from USD 279.22 billion in 2024 to USD 1,811.75 billion by 2030 (Grand View Research, Inc., 2025). This momentum highlights the need to better understand how PTMs are being reused in practice.
@@
Integrating PTMs as dependencies has lots of problems (Jiang et al., 2024; Banyongrakkul et al., 2025). We believe that these problems signal a new form of software dependency, which we term Software Dependencies 2.0. We contrast this with Software Dependencies 1.0, by which we refer to the integration of conventional software libraries or packages, which are composed of human-written instructions (code-centric) and are often governed by mature package management ecosystems (Decan et al., 2019). Software Dependencies 2.0, by comparison, refers to the integration of software with learned behaviors encapsulated in machine learning models whose parameters are determined by data and optimization processes (model-centric). These dependencies are often distributed by model hubs with reusable artifacts, such as pre-trained parameters, architectures, and vocabularies. This shift from code-centric to model-centric dependencies may pose new challenges to the maintainability of software that incorporates such dependencies.
@@
Recent work (Jiang et al., 2024) frames PTM development and deployment as a model supply chain, spanning upstream model producers (Brown et al., 2020; Rombach et al., 2022b; OpenAI, 2023a), intermediary platforms (Castano et al., 2023; Mitchell et al., 2019), and downstream software systems that consume these models. While prior research has primarily focused on upstream concerns such as model documentation, security, and sharing (Castano et al., 2023; Jiang et al., 2022; Kathikar et al., 2023; Jiang et al., 2025), downstream reuse remains less understood. Understanding downstream PTM reuse is crucial: without it, systems may become harder to maintain and reproduce (Decan et al., 2019; Okafor et al., 2022), and upstream producers may miss key requirements for designing PTMs, platforms, and tools. From the perspective of Software Dependencies 2.0, understanding downstream practices reveals how model-centric dependencies are managed and integrated into software. By empirically examining how PTMs are reused in real-world projects, our study provides a foundation for developing tools, standards, and guidelines to support reliable and maintainable PTM integration.
@@ 2. RQ2:
    What are the stages and their organization in the pipeline of PTM-based open source projects?
@@ 3. RQ3:
To address these questions, we use the PeaTMOSS dataset (Jiang et al., 2024), which contains 28,575 GitHub repositories that employ models from Hugging Face and PyTorch Hub. From this dataset, we extract a stratified statistically representative sample of 650 popular projects (341 using PTMs from Hugging Face Hub and 314 from PyTorch Hub). After categorizing these projects by the functional role of PTMs, we focus our analysis on a manually-curated 401 projects that employ PTMs to implement their core functionalities.
@@
Combining these analyses, we report the characteristics and challenges of PTM reuse in ML-enabled systems. For RQ1, we observed diverse and ad hoc practices in structuring and declaring dependencies on PTMs (§4.1). Dependencies are often inconsistently specified across source code, configuration files, and README documents, with limited use of structured tools for version control. For RQ2, we first analyzed the stages of PTM reuse and identified ten recurring stages (§4.2). Based on how projects select and organize these stages, we observed three common types of PTM reuse pipelines: feature extraction, generative, and discriminative. Across all pipeline types, PTMs were reused either as-is or with adaptation (e.g., head addition or architectural modification). Adaptation was far more common, showing that PTM reuse is rarely plug-and-play. Compared to conventional ML pipelines (Amershi et al., 2019; Biswas et al., 2022), these PTM-centric pipelines often bypass or restructure standard stages such as model training or feature engineering. However, this shift introduces new challenges related to adapting and maintaining PTMs within downstream applications. For RQ3, we found that PTMs frequently interact with other models (in 50% of the studied projects). We developed a taxonomy of four recurring interaction types (§4.3), i.e., feature handoff, feedback guidance, evaluation, and post-processing refinement to describe cross-stage interaction between pipelines involving both PTMs and conventional ML models. These interactions, which vary from tightly coupled to modular designs, introduce additional architectural complexity that has been overlooked in previous literature (Peng et al., 2020; Sens et al., 2024).
@@
Paper organization: Section 2 discusses the background on software dependencies, PTM, and PTM reuse. Section 3 motivates the RQs and presents the study design. Section 4 details the methodology and results for RQs. Section 5 presents discussions and implications of our findings. Section 6 discusses threats to validity, Section 7 discusses the related work, and finally Section 8 concludes the article. This paper is accompanied by an artifact for reproducibility, described in Section 9.
@@ 2 Background
Reusing pre-trained models (PTMs) has become a common practice in modern software systems. Such reuse introduces a new type of software dependency that differs from conventional libraries. In this section, we provide background on three key topics. First, we review Software Dependencies. Second, we discuss PTMs as Software Dependencies 2.0, highlighting their unique characteristics. Third, we describe PTM Reuse Pipelines and cross-stage multi-model interactions, focusing on how multiple models are composed and coordinated within complex systems.
@@ 2.1 Software Dependencies
Software systems are rarely built in isolation; they rely heavily on external components that provide functionality through well-defined interfaces (Sonatype, Inc., 2024). These relationships are commonly referred to as software dependencies. A software dependency represents a relationship in which one software component relies on another to function correctly. This dependence can occur through function calls, data access, or library usage. We refer to the typical dependencies with which the reader is familiar — logging libraries, web frameworks, wrappers to interact with different services — as Software Dependencies 1.0. By this we mean the code-centric components whose behavior is explicitly programmed. In widely used package ecosystems such as PyPI and NPM, dependencies are typically declared in manifest files (e.g., requirements.txt, package.json) and resolved automatically by package managers (Wittern et al., 2016; Decan et al., 2019). In this context, dependency management primarily concerns versioning, API stability, and security risks, rather than the behavior of the dependency itself, since the functionality of these components is fully defined by human-written code (Pashchenko et al., 2020). While most dependencies require little to no configuration, there are occasional cases of components with more extensive configurability (Clements and Northrop, 2002).
@@ 2.3 PTM Reuse Pipelines and Cross-stage Multi-model Interactions
Following the discussion of PTMs and their role as dependencies in Section 2.2, we next focus on the pipelines that operationalize them. We define a PTM reuse pipeline as a structured sequence of stages through which a pretrained model progresses from acquisition to operational use. Figure 2 depicts such a pipeline and its possible interactions with other pipelines.
@@
The notion of a PTM reuse pipeline builds on the broader concept of software and data processing pipelines, introduced by Shaw and Garlan (1996) as an architectural style in which components (filters) are connected by data flows (pipes).  Garlan (2000) further developed this idea with descriptions and examples that help software developers design and communicate complex systems in a clear and intelligible manner. In conventional ML  [^0] Conventional ML refers to models trained from scratch on data., pipelines are the dominant abstraction for workflows involving data processing, modeling, training, and post-processing (Amershi et al., 2019; Biswas et al., 2022). We extend this convention by treating the PTM reuse pipeline as a reusable architectural unit distinct from both the PTM itself and the broader software system.
@@ 2. RQ2:
    What are the stages and their organization in the pipeline of PTM-based open source projects?
@@ 3. RQ3:
We begin with RQ1, which focuses on the structure and documentation of PTM dependencies. Conventional libraries are typically declared in configuration or manifest files (e.g., package.json, pom.xml) that are processed by package managers to install and manage them. PTMs, by contrast, combine executable code with learned artifacts (Jiang et al., 2023b) and are often integrated through local model checkpoints or inline hub references (Jiang et al., 2024). Currently, no standardized package manager exists for PTM dependencies, and important information about model configurations and auxiliary assets may be scattered across code, scripts, or textual files. As a result, it is difficult to trace which PTMs are reused within a project and how they are reused. This traceability forms the critical foundation for reasoning about their reuse and integration, which we explore further in RQ2 and RQ3. RQ1, therefore, takes the first step by asking: How many PTMs are reused within a project? How do these PTMs relate to one another (e.g., are they interchangeable or complementary)? And where and how are these dependencies documented? By addressing these questions, RQ1 establishes an understanding of how PTM dependencies are structured and documented in practice, providing the basis for deeper analysis in subsequent RQs.
RQ2 examines how PTM reuse pipelines diverge from conventional ML pipeline structures (see Figure 3), which typically consist of standardized stages, including data preparation, feature engineering, modeling, training, and evaluation. Because PTMs are pre-trained and reusable, PTM reuse may lead to the introduction of alternative or additional stages, such as inference, fine-tuning, or model adaptation. Characterizing these real-world reuse pipelines is essential for understanding how developers adapt and integrate PTM dependencies into their systems, revealing practical reuse patterns and the underlying complexity of integration.
@@
To address these research questions, we systematically extract and analyze PTM-related artifacts such as model loading scripts, configuration files, and downstream usage code from OSS repositories. Our analysis is further enriched with insights from hub documentation and existing integration guides. Figure 4 illustrates our methodological framework.
Figure 4: Study Overview: Projects were filtered and analyzed to investigate PTM usage, pipeline stages, and model interactions.
3.2 Data Collection and PTM Usage Tracing
@@
First, we filter by project popularity, including only repositories with at least 100 GitHub stars, an approach consistent with prior work (Han et al., 2019), to focus on impactful OSS projects. This yields 3,067 repositories that reuse models from Hugging Face and 1,707 repositories that reuse models from PyTorch Hub. Next, to ensure statistical representativeness while keeping the analysis tractable, we apply stratified random sampling by hub (Hugging Face and PyTorch Hub) with a 95% confidence level and a 5% margin of error. Stratifying by hub was chosen because the two hubs differ in their typical domains (e.g., NLP, computer vision, audio). This results in a final sample of 341 Hugging Face and 314 PyTorch Hub repositories. After removing duplicates, 650 unique projects remained.
@@
  • Extension to the Remaining Projects: The same two authors annotated the remaining projects based on the initial categories, achieving a Cohen’s kappa coefficient of 0.79 for Hugging Face and 0.80 for PyTorch, indicating substantial inter-rater agreement (Viera et al., 2005). They discussed and resolved any discrepancies.
@@
Educational Resources Projects that aim to educate developers on PTM usage, offering tutorials, documentation, and example implementations. Hugging Face Notebooks5  (research code and tutorials)
@@
These functions serve as the starting points for our tracing procedure. We then perform an iterative, breadth-first traversal of the call graph using git grep, identifying where each function is called and extracting the enclosing caller functions. We set the depth limit to 6 to balance completeness and noise. Based on a pilot study of 20 randomly sampled projects, we observed that more than 95% of PTM-related activities occur within 3 to 4 levels of indirection from the model load call. Beyond 6 levels, the likelihood of encountering unrelated or utility function references increases significantly. Since our tracing is based on static patterns, it may miss dynamic method calls (e.g., reflection, dynamic dispatch) or implicit data flows (e.g., via global state or object attributes). To mitigate these limitations, we manually inspect cases where tracing appears incomplete or ambiguous, especially in projects with non-standard control flow. For 15 projects, we have identified the relevant files during this manual inspection. This manual step helps validate results and recover connections that static analysis alone may miss.
Additional Source Collection. In addition to PTM-usage files, we identify several other types of files for further analysis. Our collection includes README files, which provide insight into the project. Configuration files are collected as they define the settings and parameters necessary for model usage, including architecture configurations, training hyperparameters, and optimization techniques. We collect all configuration files with .yaml, .yml, .xml, .config or .json extensions. [^2] In our dataset, 95% of projects that included configuration files used these extensions.

### v2

@@ 1 Introduction
Deep neural networks (DNNs) are driving major advances in domains such as image analysis (Namysl and Konya, 2019; Yin et al., 2019), natural language processing (Collobert and Weston, 2008; Nakano et al., 2021; Zhou, 2022), and speech recognition (Gaikwad et al., 2010; Abdel-Hamid et al., 2014; Nassif et al., 2019). In software engineering, DNNs are increasingly applied to tasks such as code generation and automated testing (Le et al., 2020; Ale and Yarram, 2024; Chen et al., 2025). Yet, training these models from scratch is resource-intensive, requiring substantial computational power and engineering effort as architectures grow in size and complexity (Narayanan et al., 2019; Geiping and Goldstein, 2023). Pre-trained models (PTMs) have emerged as a practical solution to this challenge. By encapsulating both the architecture and the parameters learned from data, PTMs enable developers to transfer knowledge across tasks, significantly reducing training costs and development time (Marcelino, 2018; Han et al., 2021). The recent emergence of foundation models (FMs), i.e., large-scale PTMs trained on broad, heterogeneous data and adaptable to a variety of tasks (Bommasani et al., 2021), has further accelerated PTM reuse, with models such as GPT-4 (OpenAI, 2023a) and Stable Diffusion (Rombach et al., 2022a). Reports indicate that the global AI market, encompassing technologies like PTMs, is projected to grow from USD \$279.22 billion in 2024 to USD \$1,811.75 billion by 2030 (Grand View Research, Inc., 2025). This momentum highlights the need to better understand how PTMs are being reused in practice.
@@
Integrating PTMs as dependencies introduces several technical challenges spanning model integration, environment setup, documentation, resource configuration, and adapting PTM behavior to project-specific tasks (Jiang et al., 2024; Banyongrakkul et al., 2025). We believe that these problems signal a new form of software dependency, which we term Software Dependencies 2.0. We contrast this with Software Dependencies 1.0, referring to the integration of conventional software libraries or packages built from source code (code-centric) and typically managed through mature package management ecosystems (Decan et al., 2019). Software Dependencies 2.0, by comparison, refers to the integration of software with learned behaviors encapsulated in machine learning (ML) models whose parameters are determined by data and optimization processes (model-centric). These dependencies are often distributed through model hubs as reusable artifacts and pose new challenges for the maintainability of software that incorporates them.
@@
Recent work (Jiang et al., 2024) frames PTM development and deployment as a model supply chain, spanning upstream model producers (Brown et al., 2020; Rombach et al., 2022b; OpenAI, 2023a), intermediary platforms (Castano et al., 2023; Mitchell et al., 2019), and downstream software systems that consume these models. While prior research has primarily focused on upstream concerns such as model documentation, security, and sharing (Castano et al., 2023; Jiang et al., 2022; Kathikar et al., 2023; Jiang et al., 2025), downstream reuse remains less understood. Understanding downstream PTM reuse is crucial: without it, systems may become harder to maintain and reproduce (Decan et al., 2019; Okafor et al., 2022), and upstream producers may miss key requirements for designing PTMs, platforms, and tools. From the perspective of Software Dependencies 2.0, understanding downstream practices reveals how model-centric dependencies are managed and integrated into software.
@@ 2. RQ2:
    What are the stages and their organization in the reuse pipeline of PTM-based open source projects?
@@ 3. RQ3:
To address these questions, we use the PeaTMOSS dataset (Jiang et al., 2024), which contains 28,575 GitHub repositories that employ models from Hugging Face and PyTorch Hub. From this dataset, we extract a stratified statistically representative sample of 650 popular projects (341 using PTMs from Hugging Face Hub and 314 from PyTorch Hub). After categorizing these projects by the functional role of PTMs, we focus our analysis on a manually-curated set of 401 projects that use PTMs to implement core functionality.
@@
The main findings from our RQs are:
  • RQ1: We observed various practices in structuring and declaring dependencies on PTMs (§4.1). Among multi-PTM projects, dependencies are either interchangeable (37%) or complementary (23%). Dependencies are often inconsistently specified across source code, configuration files, and README files, with limited version control.
  • RQ2: We analyzed the stages of PTM reuse and identified ten recurring stages (§4.2). Projects commonly organize these stages into three pipeline types: feature extraction, generative, and discriminative. Across pipelines, PTMs are reused either as-is or with adaptation (e.g., head addition or architectural modification), with adaptation being far more common, suggesting that PTM reuse is rarely plug-and-play. Compared to conventional ML pipelines (Amershi et al., 2019; Biswas et al., 2022), PTM-centric pipelines often bypass or restructure standard stages such as model training or feature engineering, introducing new maintenance challenges.
  • RQ3: PTMs frequently interact with other models (in 50% of the studied projects). We developed a taxonomy of four recurring interaction types (§4.3): feature handoff, feedback guidance, evaluation, and post-processing refinement.
To better capture the emerging characteristics observed in our empirical study, we introduce the concept of Software Dependencies 2.0, formally defined below.
Definition: Software Dependencies 2.0 Synthesizing the results of our empirical study, we define Software Dependencies 2.0 as model-centric dependencies whose behavior is determined by learned parameters rather than explicit source code, and whose integration requires managing semantic, pipeline-level, and cross-model interactions. This definition refines our earlier intuitive framing of Software Dependencies 2.0 as the integration of software with learned behaviors encapsulated in ML models. Unlike Software Dependencies 1.0 (code-centric), which are typically declared through static manifests and integrated via stable APIs, Software Dependencies 2.0 introduce dependencies whose meaning emerges from reuse pipelines, adaptation strategies, runtime configuration, and interactions with other models. As summarized in Table 7, these dependencies differ from conventional ones along many dimensions: Dependency structure shifts from static edges (“depends on”) to semantic and cross-stage interactions. Integration involves both code APIs and model artifacts, e.g., pretrained weights, tokenizers, and adapters, often requiring adaptation such as fine-tuning or head selection. Interchangeability and complementarity emerge at the model-family and pipeline level rather than at rigid API boundaries. Behavior is context-sensitive and probabilistic rather than deterministic. Traceability is often decentralized and implicit, i.e., scattered across code, configuration files, and runtime context, rather than centrally declared in manifests. PTM reuse provides a canonical instantiation of Software Dependencies 2.0, enabling us to ground this definition in observed practices across real-world systems.
@@
Paper organization: Section 2 discusses the background on software dependencies, PTM, and PTM reuse. Section 3 motivates the RQs and presents the study design. Section 4 details the methodology and results for RQs. Section 5 presents discussions and implications of our findings. Section 6 discusses threats to validity, Section 7 discusses related work, and Section 8 concludes. We describe our reproducibility artifact in Section LABEL:sec:declarations.
@@ 2 Background
Reusing pre-trained models (PTMs) has become a common practice in modern software systems. Such reuse introduces a new type of software dependency that differs from conventional libraries. In this section, we provide background on three key topics. First, we review Software Dependencies. Second, we discuss PTMs as “Software Dependencies 2.0”, highlighting their unique characteristics. Third, we describe PTM Reuse Pipelines and cross-stage multi-model interactions, focusing on how multiple models are composed and coordinated within complex systems.
@@ 2.1 Software Dependencies
Software systems are rarely built in isolation; they integrate external components that provide functionality through well-defined interfaces (Sonatype, Inc., 2024). These relationships are commonly referred to as software dependencies. A software dependency represents a relationship in which one software component relies on another to function correctly. This dependence can occur through function calls, data access, or library usage. We refer to the typical dependencies with which the reader is familiar — logging libraries, web frameworks, wrappers to interact with different services — as Software Dependencies 1.0. By this we mean the code-centric components whose behavior is explicitly programmed. In widely used package ecosystems such as PyPI and NPM, dependencies are typically declared in manifest files (e.g., requirements.txt, package.json) and resolved automatically by package managers (Wittern et al., 2016; Decan et al., 2019). In this context, dependency management primarily concerns versioning, API stability, and security risks, rather than the behavior of the dependency itself, since the functionality of these components is fully defined by human-written code (Pashchenko et al., 2020). While most dependencies require little to no configuration, there are occasional cases of components with more extensive configurability (Clements and Northrop, 2002).
@@ 2.3 PTM Reuse Pipelines and Cross-stage Multi-model Interactions
To characterize how PTMs are reused in practice, we introduce the notion of PTM reuse pipelines, i.e., the operational structures through which PTM dependencies are integrated, configured, and composed within software systems. A PTM reuse pipeline is a structured sequence of stages through which a pretrained model progresses from acquisition to operational use. Fig. 2 presents a conceptual framework for understanding PTM reuse pipelines, drawn from established literature on software architecture and ML workflows (Amershi et al., 2019; Biswas et al., 2022). This background framework provides the conceptual foundation for our analysis of real-world PTM usage patterns in Section 4.
@@
The notion of a PTM reuse pipeline builds on the broader concept of software and data processing pipelines, introduced by Shaw and Garlan (1996) as an architectural style in which components (filters) are connected by data flows (pipes).  Garlan (2000) further developed this idea with descriptions and examples that help software developers design and communicate complex systems in a clear and intelligible manner. In conventional ML  [^0] Conventional ML refers to models trained from scratch on data., pipelines are the dominant abstraction for workflows involving data processing, modeling, training, and post-processing (Amershi et al., 2019; Biswas et al., 2022), as shown in Fig. 3. We extend this convention by treating the PTM reuse pipeline as a reusable architectural unit distinct from both the PTM itself and the broader software system.
@@ 1. RQ1:
    How do open-source projects structure and document their PTM dependencies? To structure our analysis, we further decompose this research question into the following sub-research questions:
    • RQ1.1: How do OSS projects structure their PTM dependencies?
    • RQ1.2: Where and how are PTM dependencies documented in OSS projects?
@@ 2. RQ2:
    What are the stages and their organization in the reuse pipeline of PTM-based open source projects?To structure our analysis, we further decompose this research question into the following sub-research questions:
    • RQ2.1: Which stages characterize real-world PTM reuse pipelines, and how do they differ from conventional ML pipeline stages?
    • RQ2.2: How are these stages organized within OSS projects?
@@ 3. RQ3:
We begin with RQ1, which focuses on the structure and documentation of PTM dependencies. Conventional libraries are typically declared in configuration or manifest files (e.g., package.json, pom.xml) that are processed by package managers to install and manage them. PTMs, by contrast, combine executable code with learned artifacts (Jiang et al., 2023b) and are often integrated through local model checkpoints or inline hub references (Jiang et al., 2024). Based on prior observations, there is currently no widely adopted standardized package manager for PTM dependencies. (Jiang et al., 2024). As a result, it is difficult to trace which PTMs are reused within a project and how they are reused. This traceability forms the critical foundation for reasoning about their reuse and integration, which we explore further in RQ2 and RQ3. RQ1, therefore, takes the first step by asking: How many PTMs are reused within a project? How do these PTMs relate to one another (e.g., are they interchangeable or complementary)? And where and how are these dependencies documented? By addressing these questions, RQ1 establishes an understanding of how PTM dependencies are structured and documented in practice, providing the basis for deeper analysis in subsequent RQs.
RQ2 examines how PTM reuse pipelines diverge from conventional ML pipeline structures (see Figure 3), which typically consist of standardized stages, including data preparation, feature engineering, modeling, training, and evaluation. RQ2 focuses specifically on PTM reuse, examining how PTMs are incorporated, adapted, and operationalized within PTM reuse pipelines.
Because PTMs are pre-trained and reusable, PTM reuse may lead to the introduction of alternative or additional stages, such as inference, fine-tuning, or model adaptation. Characterizing these real-world reuse pipelines is essential for understanding how developers adapt and integrate PTM dependencies into their systems, revealing practical reuse patterns and the underlying complexity of integration.
@@
To answer these research questions, we systematically extract and analyze PTM-related artifacts such as model loading scripts, configuration files, and downstream usage code from OSS repositories. Our analysis is further enriched with insights from hub documentation and existing integration guides. Figure 4 illustrates our methodological framework.
Figure 4: Data Preparation: Projects were filtered and analyzed to investigate PTM usage, pipeline stages, and model interactions.
3.2 Data Preparation
@@
First, we filter by project popularity, including only repositories with at least 100 GitHub stars, an approach consistent with prior work (Han et al., 2019), to focus on impactful OSS projects. This yields 3,067 repositories that reuse models from Hugging Face and 1,707 repositories that reuse models from PyTorch Hub. Next, to ensure statistical representativeness while keeping the analysis tractable, we apply stratified random sampling by hub (Hugging Face and PyTorch Hub) with a 95% confidence level and a 5% margin of error. Stratifying by hub was chosen because the two hubs differ in their typical domains (e.g., NLP, computer vision, audio). This results in a final sample of 341 Hugging Face and 314 PyTorch Hub repositories. Five repositories appearing in both hubs were counted once in the total number of projects, resulting in 650 unique projects.
@@
  • Extension to the Remaining Projects: The same two authors annotated the remaining projects based on the initial categories, achieving a Cohen’s kappa coefficient of 0.79 for Hugging Face and 0.80 for PyTorch, indicating substantial inter-rater agreement (Viera et al., 2005). Disagreements were resolved through careful discussions to ensure consistent interpretation of category boundaries. Because our categorization targets high-level PTM reuse roles rather than fine-grained implementation details, the categories were intentionally coarse-grained and stabilized during the pilot phase. No additional categories emerged during the extension phase, and all disagreements were resolved through consensus.
@@
Educational resources Projects that aim to educate developers on PTM usage, offering tutorials, documentation, and example implefmentations. Hugging Face Notebooks5 (research code and tutorials)
@@
These functions serve as the starting points for our tracing procedure. We then perform an iterative, breadth-first traversal of the call graph using git grep, identifying where each function is called and extracting the enclosing caller functions. We set the depth limit to 6 to balance completeness and noise. Based on a pilot study of 120 randomly sampled projects (approximately 30% of the dataset), we observed that more than 95% of PTM-related activities occur within 3 to 4 levels of indirection from the model load call. Beyond 6 levels, the likelihood of encountering unrelated or utility function references increases significantly.
Since our tracing is based on static patterns, it may miss dynamic method calls (e.g., reflection, dynamic dispatch) or implicit data flows (e.g., via global state or object attributes). To mitigate these limitations, we manually inspect cases where tracing appears incomplete or ambiguous, especially in projects with non-standard control flow. For 15 projects, we have identified the relevant files during this manual inspection. This manual step helps validate results and recover connections that static analysis alone may miss.
Additional Source Collection. In addition to PTM-usage files, we identify several other types of files for further analysis. Our collection includes README files, which provide insight into the project. Configuration files are collected as they define the settings and parameters necessary for model usage, including architecture configurations, training hyperparameters, and optimization techniques. We collect all configuration files with .yaml.yml.xml.config, or .json extensions. [^2] In our dataset, 95% of projects that included configuration files used these extensions.
PTM Name Identification. We identify the specific PTMs used in each project by extracting model identifiers from loading statements. PeaTMOSS provides direct mappings between known loading functions and PTM names when those names are statically defined, i.e., hardcoded string literals in the source code. These mappings are already included in the dataset. For example:
  • torchvision.models.resnet18(pretrained=True) (TorchVision)
In contrast, for dynamically loaded models, names are not fixed in the code but stored in variables, passed as arguments, or read from configuration files or command-line inputs. Although PeaTMOSS includes the loading function (e.g., from_pretrained(...)), it does not identify the exact model name. We start from the files where these calls appear and manually check nearby code and other project files collected in previous steps to trace the data flow and identify the actual model names being loaded.
3.3 Statistics of Identified PTMs
In total, we identify 1,059 PTM usages across the 401 projects, involving 381 unique models. Figure 5 shows the distribution of the 381 unique models by modality (as provided by PeaTMOSS). NLP models are the most diverse (169 models, 44.4%), followed by vision (136, 35.7%), multi-modal (51, 13.4%), speech (20, 5.2%), and audio (5, 1.3%). The models span from 2012 to 2024, with the largest concentration in 2019–2022 (63.7%). Architecturally, Transformer-based models make up 61.1%, including encoder-only (e.g., BERT, ViT; 36.6%), decoder-only (e.g., GPT-2, LLaMA; 15.3%) and encoder–decoder (e.g., T5, BART; 9.2%).
Figure 5: Number of unique PTMs by modality in the target 401 projects.